Quadratic Formula:
3(x-2)^2=48

Answer by stanbon(75887)

3(x-2)^2=48
--------
(x-2)^2 = 16
---
x-2 = 4 or x-2 = -4
-------
x = 6 or x = -2
=======================
Cheers,
Stan H.
=========